A window is 50cm long. A stone is falling from a height of 40 cm above the window. It crosses the window in:

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
So the stone is falling from 40m above the windowTo find out time S=ut+1/2(gt^2)So 40= 1/2(9.8t^2) thus T= 2.85 second